OpenStudy (anonymous):

Simplify the following expression. sec^2(x) + sin^2(x) - tan^2(x) + cos^2(x) -2 0 1 2

OpenStudy (mathstudent55):

Remember the Pythagorean identities: \(\sin^2 \theta + \cos^2 \theta = 1\) \(\tan^2 \theta + 1 = \sec^2 \theta \iff \sec^2 \theta - \tan^2 \theta = 1\) Now look at what you have: \(\color{red}{sec^2 x} + \color{blue}{\sin^2 x} - \color{red}{\tan^2 x} + \color{blue}{\cos^2 x}\) Group by color and compare to the identities above.